If you weigh 200 lb force and have a mass of 5 lb mass, how much does your little sister weigh if she has a mass of 2.25 lb mass? : If you weigh 200 lb force and have a mass of 5 lb mass, how much does your little sister weigh if she has a mass of 2.25 lb mass?
Answer by stanbon(19007)

If you weigh 200 lb force and have a mass of 5 lb mass, how much does your little sister weigh if she has a mass of 2.25 lb mass?
----------
Proportion:
x/2.25 = 200/5
x/2.25 = 40
x = 2.25*40
x = 90 lb (Sister's weight)
===============
Cheers,
Stan H.